Given the function f(x)= -3x^2+ 19x - 6, algebraically determine the vertex, x-intercepts, and y-intercept.
Its vertex is: (, )
Its x-intercepts are:
Its y-intercept is:

Answer by stanbon(75887)

Given the function f(x)= -3x^2+ 19x - 6,
a = -3, b = 19
The vertex occurs at x = -b/2a = -19/(-6) = 19/6
f(19/6) = -3(19/6)^2 + 19(19/6) - 6 = 24.0833
---
Vertex(19/6 , 24.0833)
--------------------------------------------
Its x-intercepts are:
Solve -3x^2 + 19x -6 == 0
x = [-19 +- sqrt(19^2 - 4*-3*-6)]/-6
x = 1/3 or x = 6
-----------------------
Its y-intercept is:
let x = 0, then y = -6
==============================
Cheers,
Stan H.